A cylinder whose volume is 34.0 Liters contains 305 grams if elemental Oxygen at 22 degrees Celsius. How many grams of gas must be released to reduce the pressure to 1.15 atm if the temperature remains constant? Give complete solution.


1
Expert's answer
2022-02-23T11:14:05-0500

P = 1.15 atm

V = 34.0 L

T = 22C + 273.15 = 295.15 K

R = gas constant = 0.0821 L atm K−1 mol−1


Solution:

The ideal gas law can be used to determine moles of O2

PV = nRT

Rearrange the equation to isolate n:

n = PV / RT

Plug in the known values and solve for n:

n(O2) = (1.15 atm × 34.0 L) / (0.0821 L atm K−1 mol−1 × 295.15 K) = 1.6136 mol O2


The molar mass of O2 is 31.998 g/mol

Therefore,

Mass of O2 = (1.6136 mol O2) × (31.998 g O2 / 1 mol O2) = 51.63 g O2


The amount of O2 that must be released:

305 g O2 − 51.63 g O2 = 253.37 g O2


Answer: 253.37 grams of gas (O2) must be released